a ball is thrown from a point 1.02 m above the ground. the initial velocity is 18.5 m/s at an angle of 26.0 above the horizontal. find the maximum height of the ball above the ground
1
Expert's answer
2019-09-30T09:56:43-0400
"H_{max}= h_{0}+ h_{thrown}"

"h_{thrown}= \frac{V_{0}^2}{2g}\times sin^2\alpha"

Where Vo= 18.5 m/s, alpha= 26 degrees


"h_{thrown} = \frac{18.5^2}{2\times 9.81} \times 0.4384^2= 3.35 m"

"H_{max} =1.02+3.35 = 4.37 m"